OpenStudy (anonymous):

How to find derivative of sin 2/x

OpenStudy (anonymous):

Like this? \[\Large \sin \left(\frac{2}{x}\right)\]

OpenStudy (anonymous):

apply the chain rule, by multiplying the outside derivative by the inside.

OpenStudy (anonymous):

I'm take that as yes. We know that 2/x = 2x⁻¹ Now apply the chain rule: \[\frac{ d }{ dx } \sin\left(2x^{-1}\right)\]Can you handle it?

OpenStudy (anonymous):

u =2/x, cos u du/dx

OpenStudy (anonymous):

It's actually sin\[x/2\]

OpenStudy (anonymous):

Not sin 2/x It's sin x/2

OpenStudy (anonymous):

U can use the same as I showed above with x/2 instead of 2/x but it is a lot simpler, maybe you can just do it in your head without the substitution?